OpenStudy (anonymous):

Write four irrational numbers between 0.1 and 0.2.Pls explain me.

OpenStudy (anonymous):

An irrational number cannot be written as a normal fraction. For example, 1/5 is a rational number, which equals 0.2 as a decimal. But 22/7 is not, since it is repeating continuously, as 3.142857142...

OpenStudy (anonymous):

I meant π

OpenStudy (reemii):

if \(x\) is irrational, so is \(x\frac{m}{n}+p\). where m,n,p are integers. Try to find 4 different irrational numbers and by scaling and translating them you can make them fit in the interval [0.1,0.2].

OpenStudy (anonymous):

And I think it was the other way around...

OpenStudy (anonymous):

Do you know how to get the answers or do we have to show you?

OpenStudy (reemii):

example: \(\pi\) is irrational. \(\pi/10000\) is very small. then add 0.1 to it and you get a number between 0.1 and 0.2 that is irrational.

OpenStudy (reemii):

(in my earlier msg: m,n,p can be rational numbers)